# Handling Label Noise via Instance-Level Difficulty Modeling and Dynamic Optimization

### v1

@@ 4 Robust Denoising Framework
Briefly speaking, prior knowledge, i.e., wrong event information and a competitive base model is obtained in the first stage, and this prior knowledge is utilized to build on the base model with continuous noise-robust training in the second stage.
@@ 4.1 Two Stage Training
Secondly, we aim to capture a competitive base model during training to serve as the initial model for stage 2, thereby improving training quality. Recent research [34] introduced a metric called label wave, which records the number of prediction changes per epoch. The model at the first local minimum of prediction changes exhibits strong competitiveness in LNL settings. In particular, this technique does not increase the computational cost of DODO. Notably, multiple fine-tuning approaches can be employed in stage 1, including Linear Probing (LP) or Full Fine Tuning (FFT). For smaller models such as ResNet-50, we recommend using FFT to achieve a more competitive base model and improve the accuracy of wrong event estimation. For larger models such as ViT-B/16, we suggest using LP to accelerate training. The stage 1 process is illustrated in the left side of Fig.3.
Stage 2: Robust Model Adaptation. The second phase of the algorithm aims to train a noise-robust model based on the base model, incorporating prior knowledge. The overall training objective of second phase is:
$\mathcal{L}=\sum_{i=1}^{N}[\gamma_{1}(w_{i})\cdot\mathcal{L_{C}}+\epsilon(w_{i})\cdot\mathcal{L_{SIM}}+\gamma_{2}(w_{i})\cdot\mathcal{L_{N}}]$ (9)
where $\tau_{1}(\cdot)$, $\tau_{2}(\cdot)$ are the posterior probabilities of $\mathcal{B}_{1}$, $\mathcal{B}_{2}$, and $\epsilon(\cdot)$ is the difficulty coefficient in Eq. 7. We will give the formation of each loss function in detail in the following subsection. After each epoch of robust training, we also update the wrong event information for each sample. With more accurate model output, the clean distribution and the noisy distribution gradually separate, providing the model with more accurate modeling. In this way, both the model and the wrong event information become more accurate. The stage 2 process is illustrated in the right side of Fig.3.
@@ 4.2 Analysis of Loss Function
which leads to the loss term $\mathcal{L_{N}}$:
$\mathcal{L_{N}}=-f(x_{i})\ \text{log}(f(x_{i}))$ (12)
To optimize the difficult samples, we aim to enhance the feature extraction capability of the model for these samples, as their labels and model outputs are often inaccurate. Specifically, we encourage the outputs of the two views to be as close as possible, and we use $\mathcal{L_{SIM}}$ to measure this similarity:
@@
$\mathcal{L_{SIM}}$ is the mean squared error (MSE) between the two predictions. We use $\mathcal{L_{SIM}}$ to increase the loss weight for difficult samples, encouraging the model to focus more on learning from these samples and thereby achieving more robust feature extraction and classification capabilities.
@@ 5 Experiments
CoL ResNet-50 83.1 81.7 79.2 73.6 80.1 74.9 73.3 74.7
DMix ResNet-50 84.3 83.1 80.7 66.6 81.3
UNICON ResNet-50 84.1 83.2 81.1 76.8 82.1
DISC ResNet-50 83.6 82.2 80.1 77.1 82.5 75.7 74.6 75.8
DeFT CLIP-ResNet-50 84.3 82.9 79.9 64.9 79.3 76.8 73.9 74.1
@@
Table 1: Comparison with state-of-the-art LNL algorithms in test accuracy (%) on CIFAR-100 and Tiny-ImageNet datasets. For CIFAR-100, we evaluate symmetric, asymmetric, and instance noise. For Tiny-ImageNet, we evaluate symmetric and instance noise. The best results for each dataset and noise type are highlighted in bold. We report average performance of three random trials for each experiment. The results of all the baseline methods are reproduced using the open-sourced code following Implements in Section 5.1.
@@ 5.1 Experimental Settings
Synthetic Datasets. We begin by evaluating the performance of DODO on two image classification benchmarks using synthetic datasets with varying types and ratios of noisy labels. Specifically, we conducted experimental analyzes on the widely used CIFAR-100 [41] and Tiny-ImageNet [42] datasets. Given a noise transition matrix $T\in[0,1]^{K\times K}$, where $T_{ij}$ represents the probability of a ground-truth label $i$ being flipped to a corrupted label $\bar{y}=j$. Following pervious works [12, 22], we introduce three common types of label noise: 1) Symmetric Noise [6, 14]: With noise rate $\eta$ and a $K$-class image classification task, we define $T_{ij}=\frac{\eta}{K}$ for $i\neq j$, where the true labels are replaced with random labels. 2) Asymmetric Noise: $T_{ij}=p(\bar{y}=j|y=i)$, which is designed to mimic the structure of real-world label noise. The labels are only replaced by similar classes. 3) Instance-dependent Noise: $T_{ij}=p(\bar{y}=j|y=i,x)$, which represents a more realistic scenario that considers the influence of instance $x$ in the label corruption process, such as the instance-dependent noise proposed in [43].
@@
Following setting of previous work [9, 25, 8], we conduct extensive experiments on CIFAR-100 with symmetric noise ratio $r\in\{0.2,0.4,0.6\}$, asymmetric noise ratio $r\in\{0.4\}$ and instance-dependent noise ratio $r\in\{0.4\}$, and on Tiny-ImageNet with symmetric noise ratio $r\in\{0.2,0.5\}$ and instance-dependent noise ratio $r\in\{0.4\}$.
@@
We further investigate the performance of DODO on three real-world noisy label datasets: 1) CIFAR-100N [44] ($r\approx 0.4$): A variant of CIFAR-100 with real-world human annotations collected from Amazon Mechanical Turk. 2) Clothing1M [45] ($r\approx 0.385$): A large-scale dataset consisting of 1 million clothing images across 14 categories, collected from online shopping websites. 3) WebVision [1] ($r\approx 0.2$): A dataset using 1,000 classes from ImageNet ILSVRC12, containing 2.4 million images crawled from Flickr and Google. Following previous works [22, 8], we conduct experiments on the top 50 classes of the Google image subset.
@@
Architecture and baselines. Using pre-trained models for noise-robust training has been demonstrated to significantly accelerate training speed and enhance training quality [21, 46, 8, 22]. In our experiments, we primarily utilize two widely adopted pre-trained models: ViT-B/16 and ResNet-50. For ResNet-50, we compare DODO with state-of-the-art LNL algorithms, including one-stage methods SCE [47], DivideMix (DMix) [14], ELR [10], Co-learning (CoL) [25], UNICON [16], DISC [12] and two-stage methods TURN [8] and DEFT [22]. However, for ViT-B/16, we compare DODO with SCE, ELR, TURN, and DEFT, without DMix and UNICON. They train two networks simultaneously, requiring a significant number of feed-forwards for each sample, approximately 4 times and 8 times, respectively [8], limiting their scalability. To further verify the effectiveness of DODO on real-world datasets, we also make comparisons with some state-of-the-art methods including LongReMix [48], and ProMix [49].
@@
Implements. We follow the two-stage setting in TURN and DEFT. We run 5 epochs for stage one to obtain the prior knowledge about wrong event for each sample, and run 10 epochs for stage two to fully noise-resistant fine-tune the pre-trained model. For two-stage baselines, we run 5 epochs for stage one (usually obtain a clean set) and run 10 epochs for stage two to fully fine-tune the pre-trained model. For one-stage baselines, we run 15 epochs to fully fine-tune the pre-trained model. For detailed hyperparameter settings with respect to the learning rate, optimizer, and baseline, please refer to Appendix B.
@@ 5.2 Performance for Noisy Label Learning
Overall performance on synthetic datasets. Table 1 demonstrates that DODO consistently outperforms competing methods in CIFAR100 and Tiny-ImageNet with an average accuracy of 1.1%, achieving state-of-the-art performance on both large and small models. DODO outperforms DMix and UNICON especially in non-symmetric noise through its instance-level modeling and dynamic optimization. DeFT generally performs well in non-asymmetric noise, but faces challenges in resource-constrained scenarios due to its dependence on CLIP. DISC and ELR perform well in non-symmetric noise in exhibit limited correction capability for symmetric noise, which is typically the most straightforward to rectify. Each baseline has limitations in certain scenarios, while DODO, built on difficulty modeling and dynamic optimization, demonstrates strong adaptability in various types and ratios of noise.
Overall performance on real-world datasets. Table 2 demonstrates that DODO consistently outperforms competing methods in three real-world datasets with an average accuracy of 0.8%, achieving state-of-the-art performance in both large and small models. DeFT achieves promising results with dual-prompt technique but faces challenges in resource-constrained scenarios due to its dependence on CLIP. DMix and UNICON achieve competitive results in three datasets through dual-network subset partitioning, but with slower training efficiency and limitations in scalability. In challenging real-world noisy scenarios, loss regularization methods, i.e. SCE and ELR, do not perform well because they don’t attempt to modify the noisy labels.
Method Architecture CIFAR-100N Clothing1M WebVision
Standard ResNet-50 56.8 67.5 77.8
SCE ResNet-50 61.4 69.6 78.3
ELR ResNet-50 69.2 70.5 80.1
DMix ResNet-50 72.9 72.7 81.7
CoL ResNet-50 71.2 72.0 80.9
UNICON ResNet-50 73.1 72.8 82.6
DISC ResNet-50 73.3 72.3 81.3
TURN ResNet-50 69.5 68.3 79.2
DeFT CLIP-ResNet-50 65.5 70.4 76.6
@@
ELR ViT-16/B 79.0 68.5 83.9
SCE VIT-16/B 76.4 69.5 83.2
TURN ViT-16/B 78.1 70.3 83.5
ELR* CLIP-ViT-16/B 72.8 72.2 79.3
UNICON* CLIP-ViT-16/B 77.7 70.4 84.6
LongReMix* CLIP-ViT-16/B 74.0 70.6 84.9
ProMix* CLIP-ViT-16/B 76.0 70.7 84.5
DeFT* CLIP-ViT-16/B 79.1 72.5 85.2
@@
Table 2: Comparison with state-of-the-art LNL algorithms in test accuracy (%) on CIFAR-100N, Clothing1M, WebVision dataset. The best results are highlighted in bold. Results marked with an asterisk (*) are from [22], other baseline results are reproduced using the open-sourced code following Implements in Section 5.1
Figure 4: Comparison with state-of-the-art LNL algorithms in effectiveness and efficiency, utilizing pre-trained ResNet50 on the A100 40GB GPU and CIFAR-100 dataset with 20% symmetric noise.
Architecture ViT-B/16 ConvNeXt-B
Noise Sym. 60% Inst. 40% Sym. 60% Inst. 40%
Metric Acc. Time Acc. Time Acc. Time Acc. Time
DMix 90.9 1558 89.5 1596 91.0 3150 88.5 3197
DODO 91.3 443 91.9 458 90.8 779 90.8 785
Table 3: Comparison of methods scalability, running DMix and DODO with larger models ViT-B/16 and ConvNeXt-B on theA100 40GB GPU. We record the accuracy (%) and average training time (sec) per epoch.
@@ 5.3 Further Analysis
Efficiency and scalability of DODO. Existing methods have two drawbacks: high memory demand and low computation efficiency. For instance, DMix and UNICON train two networks simultaneously, which consumes large memory and limits scalability. Moreover, each sample requires 4 to 8 feed-forwards computations, resulting in low efficiency. DODO, in contrast, only requires training a single network by using wrong event metric, which can be scaled to larger models. Additionally, each sample only needs two forward computations, achieving higher efficiency. Fig.4 shows that DODO achieves state-of-the-art performance while achieving high efficiency. Table 3 shows DODO performs better on average and nearly 4 times faster than DMix on ViT-B/16 and ConvNeXt-B.
Wrong event VS. other metrics. We adjust the BMM fitting distributions of different metrics to compare their performance in noise modeling. We evaluate the modeling ability of different metrics at different training stages by changing the duration of stage 1. The results are presented in Table 4. Loss shows optimal performance in the proper early stage; however, performance rapidly degrades in later stages due to noise overfitting. Although the EMA-loss (Exponential Moving Average loss) slows the speed of noise overfitting in later phases, it concurrently restricts the model’s ability to distinguish noisy samples during the initial phase. Both forgetting event (FE) and First k-epoch Learning (FkL) maintain their ability to identify noise after the model begins to overfit the noise. However, these metrics need to accumulate over time to show noticeable differences, resulting in suboptimal performance in the early stages. Overall, wrong event exhibits superior noise modeling capabilities in all training stages. The experiment results align with our theoretical analysis in Section, confirming the exceptional efficacy of wrong event in noise identification.
@@
Epoch 1 2 4 8 16 1 2 4 8 16
Single Loss 80.2 81.1 79.5 76.3 73.2 78.2 76.9 74.6 68.8 63.2
EMA Loss 79.9 80.9 79.3 77.9 74.9 77.8 74.5 73.6 69.8 64.5
FkL - 71.8 73.2 78.2 79.6 - 72.9 75.2 80.2 80.5
FE - 65.8 71.9 75.8 78.9 - 67.0 73.2 81.7 80.8
@@
Table 4: Comparison of noise modeling ability across five metrics at four training stages in phase 1, utilizing pre-trained ResNet50 on the CIFAR100 dataset with two noise settings. The training duration for phase 2 was set to 6 epochs by default. The ’-’ notation indicates cases where BMM fails to provide adequate fitting, due to all samples being assigned an identical metric value of 0.
@@
Table 5: Ablation studies on the loss modules were conducted using ResNet50 under three different noise settings on CIFAR100.
@@
The effect of different loss term. Our loss function consists of three components: $\mathcal{L_{C}}$, $\mathcal{L_{N}}$ and $\mathcal{L_{SIM}}$, modeling clean samples, noisy samples, and difficult samples, respectively. Our ablation experiments demonstrate that all three components contribute to improving the model performance (see Table 5). We observe that $\mathcal{L_{N}}$ provides higher improvement for symmetric noise, leveraging the accuracy of model outputs to enhance classification performance. On the other hand, $\mathcal{L_{SIM}}$ shows more significant improvements for asymmetric and instance noise, improving the model ability to learn from difficult samples by enforcing consistency in predictions.
Appendix A More Empirical Observation of Wrong Event
We show a comparison between wrong event and loss under symmetric noise (see Fig.1). We also conduct experiments under asymmetric noise, and instance-dependent noise to verify that wrong event outperforms the loss across all types of noise. Fig.5 and Fig.6 show a comparison between wrong event and loss under asymmetric noise and instance noise, respectively. The experiment results reveal that in more challenging noisy environments, the model tends to rapidly fit the loss of noisy samples, resulting in compromised accuracy and stability of the provided modeling information. In contrast, the wrong event metric maintains its capability to deliver robust modeling information regarding sample cleanliness and difficulty.
Figure 5: A comparison between wrong event and loss. The baseline model is ResNet-18 trained on CIFAR-10 with 40% asymmetric noise for 100 epochs. We show the loss distributions (upper row) and wrong event distribution (lower row) during training. The four columns, (a) (b) (c) (d), represent the distributions at epoch 10, 50, 100 and during the entire training phase where the heavy lines represent the median value and the shaded areas are the interquartile ranges, respectively. It can be seen that wrong event can clearly separate noisy data and clean data, even if the model fits noise.
Figure 6: A comparison between wrong event and loss. The baseline model is ResNet-18 trained on CIFAR-10 with 40% instance noise for 100 epochs. We show the loss distributions (upper row) and wrong event distribution (lower row) during training. The four columns, (a) (b) (c) (d), represent the distributions at epoch 10, 50, 100 and during the entire training phase where the heavy lines represent the median value and the shaded areas are the interquartile ranges, respectively. It can be seen that wrong event can clearly separate noisy data and clean data, even if the model fits noise.

### v2

@@ 4 Robust denoising framework
We design a two-stage learning framework (see Fig.3), which obtains the prior knowledge, i.e., wrong event information and a competitive base model in the first stage, and then utilizes this prior knowledge to produce a noise-robust model in the second stage.
Figure 3: Illustration of the proposed IDO framework. The training process is divided into two stages. In the first stage, prior knowledge, i.e., coarse distribution of wrong event, is obtained, and a base model that owns basic discrimination capability is captured. In the second stage, robust noise learning is performed. By using BMM, we obtain both cleanliness and difficulty information for individual samples, enabling instance-level dynamic optimization. The sample’s wrong event information and the model’s classification capability mutually benefit each other, leading to high improvement.
@@ 4.1 Two stage training
Secondly, we aim to capture a competitive base model during training to serve as the initial model for stage 2, thereby improving training quality. Recent research [37] introduced a metric called label wave, which records the number of prediction changes per epoch. The model at the first local minimum of prediction changes exhibits strong competitiveness in LNL settings. In particular, this technique does not increase the computational cost of IDO. Notably, multiple fine-tuning approaches can be employed in stage 1, including Linear Probing (LP) or Full Fine Tuning (FFT). For smaller models such as ResNet-50, we recommend using FFT to achieve a more competitive base model and improve the accuracy of wrong event estimation. For larger models such as ViT-B/16, we suggest using LP to accelerate training. The stage 1 process is illustrated in the left side of Fig.3.
Stage 2: Robust Model Adaptation. The second phase of the algorithm aims to train a noise-robust model on top of the base model, incorporating the prior knowledge acquired in the first phase. The overall training objective of second phase is:
$\mathcal{L}=\sum_{i=1}^{N}[\tau_{1}(w_{i})\cdot\mathcal{L_{C}}+\epsilon(w_{i})\cdot\mathcal{L_{SIM}}+\tau_{2}(w_{i})\cdot\mathcal{L_{N}}]$ (8)
where $\tau_{1}(\cdot)$, $\tau_{2}(\cdot)$ are the posterior probabilities of $\mathcal{B}_{1}$, $\mathcal{B}_{2}$, and $\epsilon(\cdot)$ is the difficulty coefficient in Eq.5. We will give the formation of each loss function in detail in the following subsection. After each epoch of robust training, we also update the wrong event information for each sample. With more accurate model output, the clean and noisy distribution gradually separate. In this way, both the model and the wrong event information become more accurate. The stage 2 process is illustrated in the right side of Fig.3.
@@ 4.2 Analysis of loss function
which leads to the loss term $\mathcal{L_{N}}$, controlled by a confidence coefficient:
$\mathcal{L_{N}}=-c\cdot f(x_{i})\ \text{log}(f(x_{i})),c=\max(f(x_{i}))$ (11)
To optimize the difficult samples, we aim to enhance the feature extraction capability of the model w.r.t. these samples, as their labels and model outputs are often inaccurate. Specifically, we encourage the outputs of the two views to be as close as possible, and we use $\mathcal{L_{SIM}}$ to measure this similarity:
@@
$\mathcal{L_{SIM}}$ is the mean squared error (MSE) between the two predictions. We use $\mathcal{L_{SIM}}$ to increase the loss weight for difficult samples. This encourages the model to focus more on learning from these samples, thereby achieving more robust feature extraction and classification capabilities.
@@ 5 Experiments
Table 1: Comparison with state-of-the-art LNL algorithms in test accuracy (%) on CIFAR-100 and Tiny-ImageNet datasets.
@@
CoL ResNet-50 83.1 81.7 79.2 73.6 80.1 70.9 67.3 72.7
DMix ResNet-50 84.3 83.1 80.7 66.6 81.3 73.6 71.4 73.4
UNICON ResNet-50 84.1 83.2 81.1 76.8 82.2 75.6 73.4 74.7
DISC ResNet-50 83.6 82.2 80.1 77.1 82.3 75.7 74.6 75.8
DeFT CLIP-ResNet-50 84.3 82.9 79.9 69.9 82.5 76.8 73.9 74.1
@@ 5.1 Experimental settings
Synthetic Datasets. We begin by evaluating the performance of IDO on three popular image classification benchmarks (CIFAR-10, CIFAR-100 [40] and Tiny-ImageNet [41]) using synthetic datasets with varying types and ratios of noisy labels. For a noise transition matrix $T\in[0,1]^{K\times K}$, $T_{ij}$ represents the probability of a ground-truth label $i$ being flipped to a corrupted label $\bar{y}=j$. Following previous works [11, 13], we introduce three common types of label noise: 1) Symmetric Noise [6, 16]: with noise rate $\eta$ and a $K$-class image classification task, we define $T_{ij}=\frac{\eta}{K}$ for $i\neq j$, where the true labels are replaced with random labels. 2) Asymmetric Noise: $T_{ij}=p(\bar{y}=j|y=i)$, which is designed to mimic the structure of real-world label noise. The labels are only replaced by similar classes. 3) Instance-dependent Noise [42]: $T_{ij}=p(\bar{y}=j|y=i,x)$, which represents a more realistic scenario that considers the influence of instance $x$ in the label corruption process.
@@
Following previous works [24, 23], we conduct experiments on CIFAR-100 with symmetric noise ratio $r\in\{0.2,0.4,0.6\}$, asymmetric noise ratio $r=0.4$ and instance noise ratio $r=0.4$, and on Tiny-ImageNet with symmetric noise ratio $r\in\{0.2,0.5\}$ and instance noise ratio $r=0.4$.
@@
Real-World Datasets. We further investigate the performance of IDO on three real-world noisy label datasets: 1) CIFAR-100N [43] ($r\approx 0.4$): A variant of CIFAR-100 with real-world human annotations collected from Amazon Mechanical Turk. 2) Clothing1M [44] ($r\approx 0.385$): A large-scale dataset consisting of 1 million clothing images across 14 categories, collected from online shopping websites. 3) WebVision [1] ($r\approx 0.2$): A dataset using 1,000 classes from ImageNet ILSVRC12, containing 2.4 million images crawled from Flickr and Google. Following previous works [13, 12], we conduct experiments on the top 50 classes of the Google image subset.
@@
Architecture and baselines. Using pre-trained models for noise-robust training can improve training quality [27, 45, 12, 13]. In our experiments, we primarily utilize two widely adopted pre-trained models: ViT-B/16 and ResNet-50. For ResNet-50, we compare IDO with state-of-the-art LNL algorithms, including one-stage methods SCE [46], DivideMix (DMix) [16], ELR [8], Co-learning (CoL) [23], UNICON [15], DISC [11] and two-stage methods TURN [12] and DEFT [13]. For ViT-B/16, we compare IDO with SCE, ELR, TURN and DEFT. Scalability analysis (see Table 4) shows DMix and UNICON require 4× more time and 3× more memory, limiting scalability and making comprehensive comparisons on ViT-B/16 impractical. To verify the effectiveness of IDO on real-world datasets, we also compare with some state-of-the-art methods including LongReMix [47] and ProMix [48]. The results of the baseline are reproduced by using the open-sourced code.
@@
Implementation. Following setting in [12, 13]. We run 5 epochs for stage one to obtain the prior knowledge about wrong event for each sample, and run 10 epochs for stage two to fully robust train the pre-trained model. For two-stage baselines, we run 5 epochs for stage one and run 10 epochs for stage two to fully train the model. For one-stage baselines, we run 15 epochs to fully train the model. All experiment results are the averages of five random runs on a single A100 80G GPU. For detailed hyperparameter settings about the optimizer, dataset and baseline, please refer to Appendix B.
@@ 5.2 Performance for noisy label learning
Overall performance on synthetic datasets. Table 1 demonstrates that IDO consistently outperforms competing methods in CIFAR100 and Tiny-ImageNet with an average accuracy of 1.4%, achieving state-of-the-art performance on both large and small models. IDO outperforms DMix and UNICON especially in non-symmetric noise through its instance-level modeling. DeFT generally performs well in non-asymmetric noise, but faces challenges in resource-constrained scenarios due to its dependence on CLIP. DISC and ELR achieve good performance in non-symmetric noise but show limitation in symmetric noise, due to inadequate correction for symmetric noise. While existing methods struggle in some noise cases, IDO is able to handle different noise levels and types more effectively by difficulty modeling and dynamic optimization. Results of CIFAR-10 are in Appendix E.4.
Overall performance on real-world datasets. Table 2 demonstrates that IDO consistently outperforms competing methods in three real-world datasets with an average accuracy of 0.7% on both large and small models. DeFT achieves promising results with dual-prompt technique but faces challenges in fine-grained classification tasks, such as Clothing1M. DMix and UNICON achieve competitive results in three datasets through dual-network subset partitioning, but with slower training efficiency and limitations in scalability. In challenging real-world noisy scenarios, loss regularization methods, i.e. SCE and ELR, do not perform well because they do not attempt to modify the noisy labels.
Table 2: Comparison with state-of-the-art LNL algorithms in test accuracy (%) on CIFAR-100N, Clothing1M, WebVision dataset. The best results are highlighted in bold. Results marked with an asterisk (*) are from [13].
Method Architecture CIFAR-100N Clothing1M-F Clothing1M-R WebVision
Standard ResNet-50 56.8 67.5 69.8 77.8
SCE ResNet-50 61.4 69.6 71.8 78.3
TURN ResNet-50 69.5 68.3 72.9 79.2
DeFT CLIP-ResNet-50 65.5 70.4 73.3 76.6
ELR ResNet-50 69.2 70.5 73.1 80.1
CoL ResNet-50 71.2 72.0 73.0 80.9
DMix ResNet-50 72.9 72.2 74.5 81.7
UNICON ResNet-50 73.1 72.4 74.7 82.5
DISC ResNet-50 73.3 72.1 74.1 82.7
@@
SCE VIT-16/B 76.4 69.5 73.2 83.2
TURN ViT-16/B 78.1 70.3 73.9 83.5
ELR ViT-16/B 79.0 72.2 74.3 83.9
UNICON* CLIP-ViT-16/B 77.7 70.4 - 84.6
LongReMix* CLIP-ViT-16/B 74.0 70.6 - 84.9
ProMix* CLIP-ViT-16/B 76.0 70.7 - 84.5
DeFT* CLIP-ViT-16/B 79.1 72.5 - 85.2
@@ 5.3 Further analysis
Dynamic loss VS. hyperparameter-based methods. Existing methods introduce extra hyperparameters, such as coefficients of loss terms, a cutoff threshold for grouping the samples and momentum coefficient, which are hard to tune. To demonstrate the advantages of dynamic loss, we conducted experiments comparing IDO with two classic algorithms, ELR and DMix. We adjusted the hyperparameters of them to compare with IDO. The results in Table 6 and Table 7 indicate that the value of the hyperparameters for ELR and DMix greatly affects the performance. Besides, the optimal parameters vary across different noise types and ratios, while IDO achieves the best results without hyperparameter tuning. For detailed information, please refer to Appendix A.
Efficiency and scalability of IDO. Existing methods have two drawbacks: high memory demand and low computing efficiency. For instance, DMix and UNICON train two networks simultaneously, which consumes a large amount of memory and limits scalability. Moreover, each sample requires 4 to 8 feed-forward computations, resulting in low efficiency. IDO, in contrast, only requires training a single network using wrong event metric, which can be scaled to larger models. Additionally, each sample only needs two forward computations (two-views), achieving higher efficiency. Fig.4 shows that IDO achieves state-of-the-art performance while achieving high efficiency. Table 4 shows IDO is 4$\times$ faster and 3$\times$ less memory than DMix and UNICON on large models, with better performance on average. For detailed training time analysis, please refer to Appendix D.
Figure 4: Comparison with state-of-the-art LNL algorithms in effectiveness and efficiency, using pre-trained ResNet50 on the A100 80GB GPU and CIFAR-100 dataset with 20% symmetric noise.
Wrong event VS. other metrics. We adjust the BMM fitting distributions of different metrics to compare their performance in noise modeling. We change the duration of stage 1 to evaluate the metrics at different training stages. The results are presented in Table 4. Loss shows optimal performance in the early stage, but the optimal point is hard to identify; Even worse, the performance rapidly degrades in later stages due to noise overfitting. While the EMA-loss (Exponential Moving Average) mitigates noise overfitting in later training phases, it limits the capacity during the initial phase, creating a trade-off between early-stage and late-stage. The visualization about the clean sample selection ability of loss and wrong event is in Appendix C.4. Both forgetting event (FE) and First k-epoch Learning (FkL) show their noise modeling capabilities after the model begins to overfit the noise. However, these metrics need to accumulate over time to show differences, resulting in suboptimal performance in the early stages. Overall, wrong event exhibits superior noise modeling capabilities in all training stages. The experiment results match our theoretical analysis in Section 3.2, confirming the effectiveness of wrong event in identifying noise. More analysis on metrics is in Appendix C.2 C.3.
Table 3: Comparison of methods scalability, running DMix, UNICON and IDO on CIFAR-100 Inst. 40% noise with larger models on the A100 80GB GPU with a batch size of 64. We record the accuracy, average per epoch time (s) and the maximum GPU memory consumption (GB).
Model ViT-B/16 ConvNeXt-B
Metric Acc. Time Mem. Acc. Time Mem.
DMix 89.5 1596 31.8 88.5 3197 57.4
UNICON 90.5 2628 45.2 89.9 4483 78.9
[…]
Table 4: Comparison of noise modeling ability across 5 metrics at 4 training time in phase 1, using pre-trained ResNet50 on CIFAR-100. The training duration for phase 2 is set to 6 epochs. The ’-’ notation indicates cases where BMM fails to provide adequate fitting, due to all samples being assigned an identical metric value of 0.
@@
Epoch 1 2 4 8 1 2 4 8
Single Loss 80.2 81.1 79.5 76.3 78.2 76.9 74.6 68.8
EMA Loss 79.9 80.9 79.3 77.9 77.8 74.5 73.6 69.8
FkL - 71.8 73.2 78.2 - 72.9 75.2 80.2
FE - 65.8 71.9 75.8 - 67.0 73.2 81.7
@@
Table 5: Ablation studies on the loss modules are conducted using ResNet50 under three different noise settings on CIFAR100.
@@
The effect of different loss terms. Our loss function consists of three components: $\mathcal{L_{C}}$, $\mathcal{L_{N}}$ and $\mathcal{L_{SIM}}$, modeling clean samples, noisy samples, and difficult samples, respectively. Our ablation experiments demonstrate that all three components contribute to improving the model performance (see Table 5). We observe that $\mathcal{L_{N}}$ achieves a higher improvement for symmetric noise, leveraging the accuracy of model output to improve classification performance, while $\mathcal{L_{SIM}}$ shows greater improvements for asymmetric and instance noise, improving the model’s ability to learn from difficult samples by imposing consistency in predictions. More exploratory experiments are in Appendix E.
Appendix A Sensitivity analysis of parameter configurations in two SOTA methods
A.1 Analysis and experiments of ELR
ELR algorithm incorporates two crucial hyperparameters: the regularization term weight $\lambda$ and the temporal ensembling momentum coefficient $\beta$. Table 6 demonstrates the parameter sensitivity of ELR, showing that the performance is significantly influenced by parameter choices and comparison with the result of IDO. The parameter settings recommended for training from scratch in the original paper, where the authors recommend $\beta=0.7,\lambda=3$ for symmetric noise on CIFAR100 and $\beta=0.9,\lambda=7$ for asymmetric noise on CIFAR-100. However, parameter settings undergo significant shifts in pretrained settings. Optimal performance is achieved when $\beta\in[0.3,0.5],\lambda\in[3,5]$, demonstrating that optimal hyperparameters vary across different models and datasets, necessitating careful parameter tuning. Our proposed method IDO effectively addresses this challenge through dynamically weighted optimization.
Table 6: Hyperparameter sensitivity of ELR, utilizing pre-trained ResNet50 on the CIFAR100 dataset with three noise settings. The default values are set as $\lambda=3$ and $\beta=0.7$. Each time, one parameter is perturbed while others are set to default. We consider the value ranges of $\lambda\in\{0,1,3,5,7,10\}$ and $\beta\in\{0.1,0.3,0.5,0.7,0.9,0.99\}$, following original paper [8].
Hyperparameter regularization term weight $\lambda$ momentum coefficient $\beta$ IDO
Noise Dataset 0 1 3 5 7 10 0.1 0.3 0.5 0.7 0.9 0.99
CIFAR-100 Sym. 60% 41.1 41.7 73.9 75.9 46.4 1.0 12.2 75.9 77.1 73.9 51.3 42.9 81.0
A.2 Analysis and experiments of DivideMix
DivideMix algorithm incorporates three crucial hyperparameters: the unsupervised loss term weight $\lambda_{u}$, GMM threshold $\tau$ for subset partitioning and beta distribution parameter $\alpha$ of Mixup [49]. Table 7 demonstrates the parameter sensitivity of DivideMix, showing that the performance is significantly influenced by parameter choices and comparison with IDO’s result. The parameter settings recommended for training from scratch in the original paper, where the authors recommend $\lambda_{u}=25,\tau\in[0.5,0.6],\alpha=4$ for symmetric noise on CIFAR100 and $\lambda_{u}=0,\tau=0.5,\alpha=4$ for asymmetric noise on CIFAR-100. However, parameter settings undergo significant shifts in pretrained settings. Optimal performance is achieved when $\lambda_{u}\in[50,150],\tau\in[0.1,0.5]$ shows that optimal hyperparameters vary between different models and data sets, which requires careful parameter tuning. Our proposed method IDO effectively addresses this challenge through dynamically weighted optimization.
Table 7: Hyperparameter sensitivity of DivideMix, utilizing pre-trained ResNet50 on the CIFAR100 dataset with three noise settings. The default values are set as $\lambda_{u}=50$, $\tau=0.5$ and $\alpha=4$. Each time, one parameter is perturbed while others are set to default. We consider the value ranges of $\lambda_{u}\in\{0,25,50,150\}$, $\tau\in\{0.1,0.3,0.5,0.7,0.9\}$ and $\alpha\in\{0.5,1,2,4\}$ following original paper [16].
Hyperparameter $\mathcal{L_{U}}$ term weight $\lambda_{u}$ GMM threshold $\tau$ Mixup parameter $\alpha$ IDO
Noise Dataset 0 25 50 150 0.1 0.3 0.5 0.7 0.9 0.5 1 2 4
CIFAR-100 Sym. 60% 80.7 79.3 79.9 80.8 81.1 80.5 79.6 67.4 29.8 80.9 79.7 79.2 78.0 81.0